OpenStudy (anonymous):

Write the equation of any straight line that passes through the origin. Use x as the input variable and y as the output variable

OpenStudy (anonymous):

y = mx where m is any real numbers? dunno wild guess

OpenStudy (anonymous):

i think y=mx also

OpenStudy (anonymous):

y = mx + b since it crosses the origin which is (0,0) so you plug in the equation 0 = 0m + b b = 0 so your equation => y = mx + 0